Question 2

  1. i)    Assume that all of the banks in Tilaknesia hold a reserve ratio of 8%.  Calculate the simple money multiplier.   
  1. Suppose that on a given day customers deposit $1,250 into their banks. Based on the simple money multiplier calculated in part i), calculate the total amount that the money supply in the banking system will eventually increase to.

Now calculate the total amount that the money supply in the banking system will eventually increase to if the reserve ratio increases to 12.5%.  Assume the amount of funds initially deposited by customers remain at $1,250.           
Answer #1

a) i. Money multiplier = 1/Reserve ratio = 1/0.08 = 12.5

ii. Total amount of money supply created = Deposits x Multiplier = 1250 x 12.5 = $ 15,625

iii) Total amount of money supply created = 1250 x 1/0.125 = 1250 x 8 = 10,000